17. If 2x, x + 10, 3x + 2 are in A.P., then x is equal to
(a) 0
(b) 2
(c) 4
(d) 6

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: d Explaination:Reason: Since 2x, x + 10 and 3x + 2 are in A.P. ∴ 2(x + 10) = 2x + (3x + 2) ⇒ 2x + 20 – 5x + 2 ⇒ 2x – 5x = 2 – 20 ⇒ 3x = 18 ⇒ x = 6